I need help putting this word problem into algabreic formula. "7 years ago I was 11 times older than my daughter, now I am 4 times older than her, how old am I?
Answer by nerdybill(7384)

"7 years ago I was 11 times older than my daughter, now I am 4 times older than her, how old am I?
.
Let x = your age
and d = your daughter's age
then
from "7 years ago I was 11 times older than my daughter"
x-7 = 11(d-7) (equation 1)
.
and from "I am 4 times older than her"
x = 4d (equation 2)
.
Solving equation 2 for d:
d = x/4
.
Substitute the above into equation 1:
x-7 = 11(d-7)
x-7 = 11(x/4-7)
Multiply both sides by 4:
4x-28 = 11(x-28)
4x-28 = 11x-308
-28 = 7x-308
280 = 7X
40 years old = x